# Accurate INT8 Training Through Dynamic Block-Level Fallback

### v2

@@ Abstract
Transformer models have achieved remarkable success across various AI applications but face significant training costs. Low-bit training, such as INT8 training, can leverage computational units with higher throughput, and has already demonstrated its effectiveness on GPT2 models with block-level quantization. However, it struggles with modern Transformer variants incorporating GLU units. This is because those variants demonstrate complex distributions of activation outliers. To address the challenge, we propose Fallback Quantization, implementing mixed-precision GEMM that dynamically falls back 8-bit to 16-bit for activation blocks containing outliers. Experiments show that our approach is robustly competent in both fine-tuning and pretraining settings. Moreover, our method achieves a 1.57x end-to-end training speedup on RTX4090 GPUs.
@@ 1 Introduction
Recently, large-scale models based on the Transformer architecture have achieved remarkable success in natural language processing and computer vision. Models such as GPT-4 (Achiam et al., 2023), Llama (Touvron et al., 2023), and ViT (Dosovitskiy, 2020) have demonstrated state-of-the-art performance across various tasks. However, as both model parameter counts and data sizes continue to increase, training these large-scale models incurs significant computational costs. To accelerate the training process and reduce costs, researchers widely adopt low-precision numerical formats such as FP16 and BF16 for optimizing matrix multiplication operations (Micikevicius et al., 2017).
Recent work has shown that fully quantized training (FQT) with even lower precision data formats such as INT8 (Xi et al., 2024) and FP8 (Peng et al., 2023) are applied in Transformer training with promising results. FP8, first introduced in NVIDIA’s Hopper (NVIDIA, 2022b) architecture, offers a large dynamic range for training Transformer models, but its support is limited to a few specific hardware. INT8 has wider support across different platforms and is sometimes faster than FP8 operations. [^1] For example, on RTX 4090, peak INT8 is 660Tops, which is 2x faster than the 330Tflops peak FP8 compute (NVIDIA, 2022a). Despite the potential efficiency and hardware compatibility advantages, INT8 training is not yet as mature as FP8, since the narrow dynamic range makes it unsuitable to handle outliers in training.
@@
While early INT8 training works (Banner et al., 2018; Zhu et al., 2020; Chen et al., 2020) adopted per-tensor quantization with applications on convolutional neural networks, there were some recent fine-grained quantization methods that succeeded in training transformers. Particularly, Switchback (Wortsman et al., 2023) quantized activation/gradient per-token and weight per-channel for training vision transformers. Jetfire (Xi et al., 2024) proposed a more accurate per-block quantization method, where each $32\times 32$ block in activation/weight/gradient matrices had a separate scale. However, the small group size of Jetfire leads to large overhead in dequantization and accumulation, making the actual speedup unsatisfactory (only reaching 39% peak flops on RTX 4090). Moreover, Jetfire was only tested on GPT-2-style models (Radford et al., 2019), while recent architectures with GLU units and enlarged dataset (Touvron et al., 2023) could be significantly harder to train.
@@
In this work, we propose an accurate and efficient INT8 training method based on block-level mix-precision quantization. We observe that the activation outlier pattern in modern GLU-based architectures is sparse, and can be covered with a small fraction of square blocks. Therefore, we propose a dynamic block fallback quantization method to allocate higher bit-width to outliers when they are detected within specific quantization blocks. Compared with existing fine-grained quantization methods, this not only isolates the impact of outliers on the entire matrix but also improves the precision of quantization block containing outliers themselves, which could carry critical information. Importantly, such mix-precision matrix multiplication (GEMM) can be implemented efficiently, similar to regular GEMMs, with minimal modifications. To further reduce memory consumption, we also integrate our training system with activation compression (Chen et al., 2021), which stores fine-grain-quantized activations for backward calculation.
@@
First, on the algorithm side, our INT8 training method successfully solves a set of challenging finetuning and pretraining tasks on the strong Llama-3.1 and Qwen-2.5 models, with lossless accuracy and overlapping training curves with BF16 baselines. This is the first time that an INT8 training method can solve such hard problems. Second, on the kernel implementation side, our mix-precision GEMM kernel reached 425 Tops on an RTX 4090, which is 2.58% faster than BF16 and 1.65% faster than Jetfire. Our framework achieves up to 1.57x end-to-end speedup and 38% activation context memory reduction compared to BF16 training.
@@ 2 Related Work
Post Training Quantization (PTQ)  (Frantar et al., 2022; Yao et al., 2022; Xiao et al., 2023; Ashkboos et al., 2024; Zhang et al., 2025, 2024) involves calibrating and converting pre-trained models to lower precision representations. Quantization Aware Training (QAT)  (Jacob et al., 2018; Dong et al., 2020; Liu et al., 2023) integrates calibration and semi-quantization during the training process to enable models to adapt to quantization errors. Both methods focus on inference-time acceleration but retain high-precision formats during training. Since they only focus on forward propagation with fixed model parameters, there exists more room for optimization.
@@
Precision Fallback is another strategy to reduce quantization errors by keeping certain data and computations at higher bit-width. Conventional mixed precision training maintains FP32 for non-matrix multiplication operations (Micikevicius et al., 2017; Peng et al., 2023), but this approach is limited to the operator level. Recent works, such as LLM.int8() (Dettmers et al., 2022; Chen et al., 2024), preserve outliers in transformer activations at FP16 precision and utilize different TensorCores through channel shuffling. However, this approach is not suitable for dynamic training processes. Xi et al. (2023) randomly preserves certain matrix computation elements at high precision to reduce errors, but fails to address the outliers in activation. Our block fallback method only performs fallback on activation blocks containing outliers, and is capable of utilizing only low-precision matrix multiplication units.
@@ 4.3 Block Fallback Quantization
10:       for $k=0$ to $\lceil K/K_{g}\rceil$ 1 on-chip do

### v3

@@ Abstract
Transformer models have achieved remarkable success across various AI applications but face significant training costs. Low-bit training, such as INT8 training, can leverage computational units with higher throughput, and has already demonstrated its effectiveness on GPT2 models with block-level quantization. However, it struggles with modern Transformer variants incorporating GLU units. This is because those variants demonstrate complex distributions of activation outliers. To address the challenge, we propose Fallback Quantization, implementing mixed-precision GEMM that dynamically falls back 8-bit to 16-bit for activation blocks containing outliers. Experiments show that our approach is robustly competent in both fine-tuning and pretraining settings. Moreover, our method achieves a 1.57$\times$ end-to-end training speedup on RTX4090 GPU.
@@ 1 Introduction
Recently, large-scale models based on the Transformer architecture have achieved remarkable success in natural language processing and computer vision. Models such as GPT-4 (Achiam et al., 2023), Llama (Touvron et al., 2023), and ViT (Dosovitskiy, 2020) have demonstrated state-of-the-art performance across various tasks. However, as both model parameter counts and data sizes continue to increase, training these large-scale models incurs significant computational costs. To accelerate the training process and reduce costs, researchers widely adopt low-precision numerical formats such as FP16 and BF16 for optimizing matrix multiplication operations (Micikevicius et al., 2017). Furthermore, recent work has shown that fully quantized training (FQT) with even lower precision data formats, such as INT8 (Xi et al., 2024) and FP8 (Peng et al., 2023), is applied in Transformer training with promising results. FP8, first introduced in NVIDIA’s Hopper (NVIDIA, 2022b) architecture, offers a large dynamic range for training Transformer models, but its support is limited to a few specific hardware. INT8 has wider support across different platforms and is sometimes faster than FP8 operations. [^1] For example, on RTX 4090, peak INT8 is 660Tops, which is 2x faster than the 330Tflops peak FP8 compute (NVIDIA, 2022a). Despite the potential efficiency and hardware compatibility advantages, INT8 training is not yet as mature as FP8, since the narrow dynamic range makes it unsuitable to handle outliers in training. While early INT8 training works (Banner et al., 2018; Zhu et al., 2020; Chen et al., 2020) adopted per-tensor quantization with applications on convolutional neural networks, there were some recent fine-grained quantization methods that succeeded in training transformers. Particularly, Switchback (Wortsman et al., 2023) quantized activation/gradient per-token and weight per-channel for training vision transformers. Jetfire (Xi et al., 2024) proposed a more accurate per-block quantization method, where each $32\times 32$ block in activation/weight/gradient matrices had a separate scale. However, the small group size of Jetfire leads to large overhead in dequantization and accumulation, making the actual speedup unsatisfactory (only reaching 39% peak flops on RTX 4090). Moreover, Jetfire was only tested on GPT-2-style models (Radford et al., 2019), while recent architectures with GLU units and enlarged dataset (Touvron et al., 2023) could be significantly harder to train.
@@
Motivation and our method. In this work, we propose an accurate and efficient INT8 training method based on block-level mix-precision quantization. We observe that the activation outlier pattern in modern GLU-based architectures is sparse, and can be covered with a small fraction of square blocks. Therefore, we propose a dynamic block fallback quantization method to allocate a higher bit-width to outliers when they are detected within specific quantization blocks. Compared with existing fine-grained quantization methods, this not only isolates the impact of outliers on the entire matrix but also improves the precision of the quantization block containing outliers themselves, which could carry critical information. Importantly, such mix-precision matrix multiplication (GEMM) can be implemented efficiently, similar to regular GEMMs, with minimal modifications. To further reduce memory consumption, we also integrate our training system with activation compression (Chen et al., 2021), which stores fine-grain-quantized activations for backward calculation.
@@
Result and contribution. First, on the algorithm side, our INT8 training method successfully solves a set of challenging finetuning and pretraining tasks on the strong Llama-3.1 and Qwen-2.5 models, with lossless accuracy and overlapping training curves with BF16 baselines. This is the first time that an INT8 training method can solve such hard problems. Second, on the kernel implementation side, our mix-precision GEMM kernel reached 425 TOPS on an RTX 4090, which is 2.58$\times$ faster than BF16 and 1.65$\times$ faster than Jetfire. Our framework achieves up to 1.57$\times$ end-to-end speedup and 38% activation context memory reduction compared to BF16 training.
@@ 2 Related Work
Post Training Quantization (PTQ)  (Frantar et al., 2022; Yao et al., 2022; Xiao et al., 2023; Ashkboos et al., 2024; Zhang et al., 2025c, a, b, d, e; Hu et al., 2025) involves calibrating and converting pre-trained models to lower precision representations. Quantization Aware Training (QAT)  (Jacob et al., 2018; Dong et al., 2020; Liu et al., 2023) integrates calibration and semi-quantization during the training process to enable models to adapt to quantization errors. Both methods focus on inference-time acceleration but retain high-precision formats during training. Since they only focus on forward propagation with fixed model parameters, there exists more room for optimization.
@@
Precision Fallback is another strategy to reduce quantization errors by keeping certain data and computations at a higher bit-width. Conventional mixed precision training maintains FP32 for non-matrix multiplication operations (Micikevicius et al., 2017; Peng et al., 2023), but this approach is limited to the operator level. Recent works, such as LLM.int8() (Dettmers et al., 2022; Chen et al., 2024), preserve outliers in transformer activations at FP16 precision and utilize different TensorCores through channel shuffling. However, this approach is not suitable for dynamic training processes. Xi et al. (2023) randomly preserves certain matrix computation elements at high precision to reduce errors, but fails to address the outliers in activation. Our block fallback method only performs fallback on activation blocks containing outliers, and is capable of utilizing only low-precision matrix multiplication units.
@@ 4.3 Block Fallback Quantization
10:       for $k=0$ to $\lceil K/K_{g}\rceil-1$ on-chip do